Please how do I solve the following problem: Three men do a piece of work, the first does 7/15, the second does 5/6 of the remainder while the third does the rest. If the rest done by the third person equal 24units, find the whole piece of work.
Thank you so much

Answer by mananth(16946)

Three men do a piece of work, the first does 7/15, the second does 5/6 of the remainder while the third does the rest. If the rest done by the third person equal 24units, find the whole piece of work.
let total work be 1
1-7/15= 8/15
I............ 7/15
1-7/15= 8/15 (balance)
5/6 of balance
II............5/6 * 8/15= 4/9
Remaining III does
III......1-(7/15+4/9)
1-(21+20)/45
1-41/45
4/45
III.............4/45
4/45 ......................24 units
1.....................?
6*45= 270 units